# Dual-VQE: A quantum algorithm to lower bound the ground-state energy

### v1

@@ Abstract
Results of numerical simulations—We now present the results of simulations that test the performance of dual-VQE on the transverse field Ising model, described by the following Hamiltonian of $n$ qubits:
$H_{\text{TFI}}=\sum_{i=1}^{n-1}J_{i}\sigma_{Z}^{i}\otimes\sigma_{Z}^{i+1}+\sum_{i=1}^{n}g_{i}\sigma_{X}^{i},$ (26)
where tensor products with identities are left implicit, $J_{i}$ is an interaction strength, and $g_{i}$ is related to the strength of the external field. In particular, we showcase our algorithm on the following two-qubit example problem instance:
$H=\sigma_{Z}\otimes\sigma_{Z}+\sigma_{X}\otimes\sigma_{I}+\sigma_{I}\otimes\sigma_{X}.$ (27)
Prior to training, we initialized the parameters $\eta$, $\mu$, and $\theta$. We set the parameters $\eta$ and $\nu$ in (12) to be equal to zero and one, respectively. We selected the components of the initial parameter vector $\theta$ for the state $\omega(\theta)$ uniformly at random from $[0,2\pi]$. The optimal state $\omega$ is almost always a full-rank mixed state. Thus, we used either a three-layer purification ansatz or a convex-combination ansatz with a two-layer quantum circuit Born machine and a two-layer unitary. The particular structure of the purification ansatz and the quantum circuit Born machine we use can be found in a figure in our companion paper [17].
@@
The results of the simulations are displayed in Figure 1. In these plots, solid lines depict results of both the standard VQE algorithm (with fixed learning rate of $0.005$ and gradient estimated by normalized SPSA) and our dual-VQE approach. The solid lines show median values across ten independent runs, and the shading is the interquartile range. The dashed line is the true value of the ground-state energy estimated via the SDP. The variation displayed in the graphs comes from three sources of randomness: the randomized initializations of all parameterized quantum circuits, randomness in the SPSA algorithm for gradient estimation, and shot noise over $10^{12}$ shots for every measurement of a quantum circuit. In our simulations, we were able to achieve an error of order $10^{-2}$ after $20,000$ iterations of training. These results give evidence that using VQE and dual-VQE allows us to bound the true value from both above and below, respectively, as desired.
@@
Discussion and conclusion—In conclusion, we proposed dual-VQE as a variational quantum algorithm for lower bounding the ground-state energy of a Hamiltonian. This lower bound can be compared with the upper bound obtained from the traditional VQE method, and the two bounds serve as quality checks on each other that ideally determine an interval containing the optimal ground-state energy. As discussed above, the quantum computational resources required for evaluating the objective function of dual-VQE are no greater than those needed for VQE’s objective function when using the convex combination ansatz in conjunction with a classical generative model.
@@
While Theorem 1 guarantees that dual-VQE provides a rigorous lower bound on the ground state energy in the limit $c\rightarrow\infty$, for finite $c$ the value provided by dual-VQE is technically an estimate rather than a strict bound. However, our numerical simulations give evidence, for the example considered, that VQE and dual-VQE do indeed bound the optimal ground-state energy from above and below, respectively. Nonetheless, it would be valuable to investigate whether it is possible to obtain bounds on the degree to which deviations are possible for finite $c$. This would allow us to certify the reliability of dual-VQE and serve as a guide to an optimum choice for $c$.
A key issue that plagues variational quantum algorithms is the presence of barren plateaus in the training landscape [31, 20, 37], meaning that the gradient (or equivalently cost differences [1]) vanishes exponentially fast in the number of qubits. This has been shown to occur for highly expressive [24] or highly entangling [32] ansätze or even at low circuit depths for global costs [15]. For local problem Hamiltonians, the only term we need to evaluate on a quantum computer, $\text{Tr}[H\omega(\theta)]$, is local and hence in such cases, for $\log(n)$-depth hardware efficient ansätze, dual-VQE provably avoids barren plateaus [15]. However, a worry here is that such short-depth circuits might also be classically simulable via state-of-the-art methods [43, 4, 25, 3, 38]. An important question for dual-VQE, shared also with other variational quantum algorithms, is thus whether we can identify quantum models that both avoid barren plateaus and cannot be classically simulated. One potential avenue would be to try and find clever initialization strategies to explore small subsections of the landscape that exhibit substantial gradients and are also hard to simulate classically.
@@
For future work, it would certainly be of interest to scale up dual-VQE to much larger systems beyond the reach of brute-force classical simulation and to compare the results of VQE and dual-VQE with each other. We also think it would be worthwhile to pursue other methods, besides the penalty method, for performing the optimization in dual-VQE. A promising approach would be to combine a variational approach with an interior-point method, but for this approach, it seems we would need an efficient method, suitable for near-term quantum computers, for estimating the logarithm of the determinant of a matrix on a quantum computer.
@@
Data availability statement—All source codes used to run the simulations and generate the figures are available as arXiv ancillary files with the arXiv posting of our companion paper [17].
@@ Acknowledgements.
We thank Paul Alsing, Ziv Goldfeld, Daniel Koch, Saahil Patel, and Manuel S. Rudolph for helpful discussions. JC and HW acknowledge support from the Engineering Learning Initiative in Cornell University’s College of Engineering. ZH acknowledges support from the Sandoz Family Foundation Monique de Meuron program for Academic Promotion. IL, TN, DP, SR, and MMW acknowledge support from the School of Electrical and Computer Engineering at Cornell University. TN, DP, SR, and MMW acknowledge support from the National Science Foundation under Grant No. 2315398. DP, SR, and MMW acknowledge support from AFRL under agreement no. FA8750-23-2-0031. This material is based on research sponsored by Air Force Research Laboratory under agreement number FA8750-23-2-0031. The U.S. Government is authorized to reproduced and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of Air Force Research Laboratory or the U.S. Government. This research was conducted with support from the Cornell University Center for Advanced Computing, which receives funding from Cornell University, the National Science Foundation, and members of its Partner Program.

### v2

@@ IV Results
IV.1 Purification and convex combination without pretraining
The results of experiments using the purification and convex combination ansätze are displayed in Figure 2. The plots shown are for $n=2$ qubits. In these plots, solid lines depict results of both the standard VQE algorithm (with fixed learning rate of $0.005$ and gradient estimated by normalized Simultaneous Perturbation Stochastic Approximation (SPSA)) and our dual-VQE approach. The solid lines show median values across ten independent runs, and the shading is the interquartile range. The dashed line is the true value of the ground-state energy estimated via the SDP.
Prior to training, we initialized the parameters $\eta$, $\mu$, and $\theta$. We set the parameters $\eta$ and $\nu$ in (13) to be equal to zero and one, respectively. We selected the components of the initial parameter vector $\theta$ for the state $\omega(\theta)$ uniformly at random from $0,2\pi]$. The optimal state $\omega$ is almost always a full-rank mixed state. Thus, we used either a three-layer purification ansatz or a convex-combination ansatz with a two-layer quantum circuit Born machine and a two-layer unitary. The particular structure of the purification ansatz and the quantum circuit Born machine we use can be found in a figure in our companion paper [8].
@@ IV.1 Purification and convex combination without pretraining
The variation displayed in the graphs comes from three sources of randomness: the randomized initializations of all parameterized quantum circuits, randomness in the SPSA algorithm for gradient estimation, and shot noise over $10^{12}$ shots for every measurement of a quantum circuit. In our simulations, we were able to achieve an error of order $10^{-2}$ after $20,000$ iterations of training. These results give evidence that using VQE and dual-VQE allows us to bound the true value from both above and below, respectively, as desired.
IV.2 MPS pretraining with purification ansatz
Figure 3: Objective function and penalty value for the Ising model problem using dual-VQE and MPS pretraining. After 2,000 iterations of pretraining followed by 25,000 iterations of quantum training, the true solution is reached when $\chi_{\max}=8$, with a relative error of 0.5%.
Figure 4: Fidelity of translation between the final MPS state after pretraining and the translated quantum circuit.
The results of the different experiments where we used MPS pretraining are given in Figure 3. The plots shown are for $n=3$ qubits. To demonstrate the effectiveness of pretraining, we used three different max bond dimension values $\chi_{\operatorname{max}}=\{2,4,8\}$ and compare it with the no-pretraining case. In each of the above cases, we conducted five runs and report the mean and standard deviation.
For the three pretraining runs, with $\chi_{\operatorname{max}}=\{2,4,8\}$, we used a randomly initialized MPS on six qubits to realize a full-rank mixed state state on three qubits. We set the initial penalty parameter to be $c=30$ and kept it constant for the pretraining phase. We used gradient descent to train for $2000$ iterations, labeled by Pretraining in Figure 3.
Following the pretraining, we translated the MPS into a quantum circuit. We used the AD decomposition to decompose the MPS into three layers of two-qubit unitaries, labelled as Disentangler in Figure 4. We plot the fidelity of translation, using the AD and OD algorithms (defined in Appendix B), in Figure 4.
The circuit we obtained then serves as the starting point for one of the OD algorithm runs. For this run, labelled as Warm Start, we used 2000 iterations and a constant learning rate of $\beta=0.2$ (see (31)). We also ran the OD algorithm on a randomly initialized three layer circuit and trained for 2000 iterations, labelled as Random Start. Finally, we picked the best translated circuit. For $\chi_{\max}=8$, the translation fidelity was around $99.5\%$.
Lastly, we added a single layer of parameterized gates to the translated circuits, as described in Section III.2.3, and all the parameters are trained using a combination of a finite difference method and the SPSA algorithm. For the parameters $\eta$ and $\nu$, we used a finite difference method with a perturbation of $10^{-6}$. For the circuit parameters, we used the SPSA algorithm since it provides an unbiased estimate of the gradient with a runtime complexity that is independent of the number of parameters. For $\eta$ and $\mu$, we initialized the learning rate, denoted by $\beta_{\eta,\mu}$, to be $10^{-3}$ and for the circuit parameters, we initialized the learning rate, denoted by $\beta_{\omega}$, to be $3\times 10^{-3}$. Every 200 iterations, we modified the learning rates and the penalty parameter $c$. For the learning rates, we calculated the slope of the last 200 iterations. If the slope was negative, we set
$\beta_{\omega}$ $=\max(0.9\beta_{\omega},\beta_{\min}),$ (34)
$\beta_{\eta,\mu}$ $=\max(0.9\beta_{\eta,\mu},3\beta_{\min}).$ (35)
On the other hand, if the slope was positive, we set
$\beta_{\omega}$ $=\min(1.05\beta_{\omega},\beta_{\max}),$ (36)
$\beta_{\eta,\mu}$ $=\min(1.02\beta_{\eta,\mu},3\beta_{max}).$ (37)
In the equations above, $\beta_{\max}=10^{-3}$ and $\beta_{\min}=10^{-4}$.
Similarly, we updated the penalty parameter by calculating the slope of the penalty of the last 200 iterations. If the slope was greater than $5\times 10^{-4}$, we updated $c$ to be $0.9c$, and if the slope was between $0.0$ and $5\times 10^{-4}$, we set $c=\min(1.04c,40)$. This latter case occurs when the penalty is stagnated, and usually occurs when the minimization has come close to a minima for the current $c$ value. So, we increased the $c$ value and enforced the constraint more strongly. These penalty updates deviate from the presentation in Section III, but we have found that they work well in practice for the problems we considered.
We ran the quantum training for 25,000 iterations. The plots show that pretraining even for a small number of iterations speeds up the quantum training significantly. Furthermore, we see that increasing the bond dimension further increases the advantage. For $\chi_{\max}=8$, we see that the quantum training reaches the true solution with a relative error of $0.5\%$.
V Discussion and conclusion
@@
In conclusion, we proposed dual-VQE as a variational quantum algorithm for lower bounding the ground-state energy of a Hamiltonian. This lower bound can be compared with the upper bound obtained from the traditional VQE method, and the two bounds serve as quality checks on each other that ideally determine an interval containing the optimal ground-state energy. As discussed above, the quantum computational resources required for evaluating the objective function of dual-VQE are no greater than those needed for VQE’s objective function when using the convex combination ansatz in conjunction with a classical generative model.
@@
While Theorem 1 guarantees that dual-VQE provides a rigorous lower bound on the ground-state energy in the limit $c\rightarrow\infty$, for finite $c$, the value provided by dual-VQE is technically an estimate rather than a strict bound. However, our numerical simulations give evidence, for the examples considered, that VQE and dual-VQE do indeed bound the optimal ground-state energy from above and below, respectively. Nonetheless, it would be valuable to investigate whether it is possible to obtain bounds on the degree to which deviations are possible for finite $c$. This would allow us to certify the reliability of dual-VQE and serve as a guide to an optimum choice for $c$.
@@
For future work, it would certainly be of interest to scale up dual-VQE to much larger systems beyond the reach of brute-force classical simulation and to compare the results of VQE and dual-VQE with each other. We also think it would be worthwhile to pursue other methods, besides the penalty method, for performing the optimization in dual-VQE. A promising approach would be to combine a variational approach with an interior-point method, but for this approach, it seems we would need an efficient method, suitable for near-term quantum computers, for estimating the logarithm of the determinant of a matrix on a quantum computer. The quantum algorithms put forward in [50, 51, 52] could be helpful for this purpose.
Finally, matrix product pretrainers are just one class of pretrainers that can help find advantageous initial points. Exploring other pretraining methods, including classical neural networks, may lead to more efficient optimization schemes.
@@
Data availability statement—All source codes used to run the simulations and generate the figures are available on Zenodo [53].
@@ Acknowledgements.
We thank Paul Alsing, Ziv Goldfeld, Daniel Koch, Saahil Patel, and Manuel S. Rudolph for helpful discussions. JC and HW acknowledge support from the Engineering Learning Initiative in Cornell University’s College of Engineering. ZH acknowledges support from the Sandoz Family Foundation Monique de Meuron program for Academic Promotion. IL, TN, DP, SR, KW, and MMW acknowledge support from the School of Electrical and Computer Engineering at Cornell University. TN, DP, SR, and MMW acknowledge support from the National Science Foundation under Grant No. 2315398. DP, SR, and MMW acknowledge support from AFRL under agreement no. FA8750-23-2-0031. This material is based on research sponsored by Air Force Research Laboratory under agreement number FA8750-23-2-0031. The U.S. Government is authorized to reproduced and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of Air Force Research Laboratory or the U.S. Government. This research was conducted with support from the Cornell University Center for Advanced Computing, which receives funding from Cornell University, the National Science Foundation, and members of its Partner Program.
Appendix A Matrix product states
An $n$-qubit pure state is a vector that belongs to a complex Hilbert space of dimension $2^{n}$, thus requiring $2^{n}$ complex numbers to fully specify it. Furthermore, unitary operations acting on this pure state are, in general, of dimension $2^{n}\times 2^{n}$ and thus require both an exponential amount of memory and computation to implement. In tensor-network notation, such a vector is generally represented by a sphere with a single leg to specify the index [54].
For example, an $n$-qubit pure state $|\psi\rangle$ can be expanded as
$|\psi\rangle=\sum_{i=0}^{2^{n}-1}\psi_{i}|i\rangle,$ (38)
where $\psi_{i}\in\mathbb{C}$ for all $i\in\{0,\ldots,2^{n}-1\}$, subject to $\sum_{i=0}^{2^{n}-1}\left|\psi_{i}\right|^{2}=1$, and the corresponding figure is given in Figure 5a). Here, $\psi$ represents the rank-$1$ tensor that contains the components of $|\psi\rangle$.
Figure 5: (a) Representation of a pure state in tensor-network notation. (b) Alternate representation of a pure state in tensor-network notation.
An alternate way of representing the same state involves using an $n$-legged tensor with each leg representing a single-qubit system. More concretely, the $n$-qubit pure state $|\psi\rangle$ can be expanded as
$|\psi\rangle=\sum_{i_{1},i_{2},\ldots,i_{n}=0}^{1}(A^{\psi})_{i_{1}i_{2}\cdots i_{n}}|i_{1}i_{2}\cdots i_{n}\rangle,$ (39)
and the corresponding figure is given in Figure 5b). Here, $A^{\psi}$ represents the rank-$n$ tensor that contains the components of $|\psi\rangle$, using the following mapping:
$(A^{\psi})_{i_{1}i_{2}\cdots i_{n}}=\psi_{i},$ (40)
where $i_{1}i_{2}\cdots i_{n}$ is the binary representation of $i$. Both representations above are equivalent to each other, and they represent no approximations or simplifications.
To manage the computational and spatial complexity of these large tensors, they can be factorized into a network of lower-order tensors. This factorization enables efficient storage and computation, depending on the resulting axis dimensions. One such factorization of this large tensor is in the form of a matrix product state.
Figure 6: MPS representation of the state $|\psi\rangle$.
Matrix product states have a linear architecture, with each tensor being connected to their nearest neighbours. For example, the tensor $A^{\psi}$ can be factorized into a linear chain of smaller tensors, as seen in Figure 6.
Figure 7: Analytical Decomposition algorithm. Reproduced from [46]. The steps are described in Algorithm 1.
Mathematically, this is expanded as
$|\psi\rangle$ $=(A^{\psi})_{i_{1}i_{2}\cdots i_{n}}|i_{1}i_{2}\cdots i_{n}\rangle$
$=(A^{\psi}_{1})^{b_{1}}_{i_{1}}(A^{\psi}_{2})^{b_{2}}_{b_{1}i_{2}}\cdots(A^{\psi}_{n})^{b_{n-1}}_{i_{n}}|i_{1}i_{2}\cdots i_{n}\rangle,$ (41)
where we use the Einstein repeated index summation convention. The indices labelled by $b_{k}$ are called bonds, and their dimension is called the bond dimension of the MPS. Note that, in general, the bonds can be of different dimension, in which case, the largest bond is quoted as the bond dimension of the MPS and is denoted by $\chi$. In addition to the bond dimension, the size of the output indices is called the physical dimension. In this work, since we are working with qubits, we fix the physical dimension to be $2$. Thus, the number of parameters of the entire MPS is given by $2n\chi^{2}$, as opposed to $2^{n}$ for the full state vector.
Matrix product states allow us to efficiently represent general pure states with the bond dimension controlling the level of compression. To fully express the entire set of pure states, the bond dimension needs to grow exponentially in $n$. Thus, the chosen $\chi$ can be increased to explore larger sectors of the entire Hilbert space [55]. The physical interpretation of the bond dimension is that it is a measure of the entanglement present in the state. Thus, MPSs with small bond dimension $\chi$ represent states with low entanglement. When used in optimization tasks, tuning the bond dimension provides for a method to smoothly increase the entanglement present in an ansatz, exploring more entangled states in the Hilbert space.
The advantages of using tensor networks do not end here. Matrix product operators generalize unitary operations and can be implemented using tensor contractions in a runtime polynomial in the bond dimension. For states with entanglement greater than the max bond dimension can exactly reproduce, the singular value decomposition may be used to find a low-bond dimension MPS approximation [56].
Just as mixed states in a Hilbert space can be represented by a pure state in a larger Hilbert space, so too can larger rank MPSs be used to represent mixed states on a smaller rank. For example, consider a mixed state $\omega_{S}$ that has a purification $|\psi\rangle_{RS}$. If $|\psi\rangle_{RS}$ has an MPS representation, then $\omega_{S}$ can be realized as an MPO using this MPS, as depicted in Figure 1a).
Figure 8: Optimizing Decomposition algorithm. Reproduced from [46]. The steps are described in Algorithm 2.
Appendix B Algorithms from RCMAP23
In this appendix, we review some algorithms studied in [46]. The algorithms are a combination of the Analytic decomposition (AD) and the Optimizing decomposition (OD) algorithms. The AD algorithm uses MPS truncation to find the optimal layer of two-qubit unitaries to disentangle the current MPS, i.e., to reduce its bond dimension. The process is repeated until a fixed iteration count is met or the fidelity reaches a threshold. On the other hand, the OD algorithm begins with a random circuit of two-qubit unitaries and systematically replaces each one with a more optimal unitary. For the sake of clarity and completeness, we provide a detailed description of both the AD and OD algorithms below.
B.0.1 Analytic Decomposition (AD)
The Analytic Decomposition (AD) algorithm was proposed as a method for decomposing an MPS of arbitrary bond dimension $\chi$ into multiple linear layers of two-qubit unitaries [45]. The AD algorithm is iterative, introducing a new layer of two-qubit unitaries in each iteration until a stopping criterion is met. The stopping criterion is usually a combination of an upper bound on the number of layers, or the fidelity between the MPS and the translated quantum circuit. These linear layers have a staircase arrangement. Let us denote such a layer of two-qubit unitaries as $L[U]$.
The algorithm is aimed at maximizing the fidelity defined by
$f(\{L[U]^{(k)}\}_{k=1}^{K})$ $=|\langle 0^{\otimes N}|\prod\limits_{k=1}^{K}{L[U]^{(k)}}^{\dagger}|\psi_{\chi_{\max}}\rangle|^{2}$
$=|\langle 0^{\otimes N}|{L[U]^{(K)}}^{\dagger}\cdots{L[U]^{(1)}}^{\dagger}|\psi_{\chi_{\max}}\rangle|^{2}$
$=|\langle\psi^{(K)}_{\text{QC}}|\psi_{\chi_{\max}}\rangle|^{2}.$ (42)
At each iteration, one layer of two-qubit unitaries is created that disentangles the truncated MPS to the all-zeros state. Then, using this layer of unitaries, the current MPS is disentangled, and then contracted, to get the MPS with lower bond dimension at the next iteration. The steps of the algorithm can be found in Algorithm 1 and a schematic can be found in Figure 7.
Algorithm 1 Analytic Decomposition (AD)
0: Input: MPS $\psi_{\chi_{\max}}$ Maximum layers $K$ Target fidelity $\hat{f}$
0:Output: Quantum Circuit layers $\Pi_{k=K}^{1}L[U]^{(k)}$
1: Set $k=1$.
2: Set $|\psi^{0}\rangle=|\psi_{\chi_{\max}}\rangle$.
3:while $k\leq K$ and $|\langle 0^{\otimes N}|\psi^{k}\rangle|^{2}<\hat{f}$ do
4:  Truncate $\psi^{k}$ to $\psi^{k}_{\chi=2}$ via SVD.
5:  Convert $\psi^{k}_{\chi=2}$ to $L[U]^{(k+1)}$.
6:  Append $L[U]^{(k+1)}$ to get $|\psi^{(k+1)}\rangle={L[U]^{(k+1)}}^{\dagger}|\psi^{(k)}\rangle$.
7:  Increment $k=k+1$.
8:end while
B.0.2 Optimizing Decomposition (OD)
The Optimizing Decomposition (OD) algorithm begins with a random layered circuit made of two-qubit unitaries and iteratively replaces each two-qubit unitary with another that increases the fidelity of the prepared state with the target MPS. The steps of the algorithm can be found in Algorithm 2 and a schematic can be found in Figure 8.
In each iteration, the environment tensor $\hat{\mathcal{F}}_{m}$ is calculated by first removing the unitary $U_{m}$ and contracting all other indices. This results in a rank-$4$ tensor, which is not necessarily unitary. It turns out that $\hat{\mathcal{F}}_{m}$ is the locally optimal operator to insert in place of $U_{m}$ to increase the fidelity. However, since it is not generally unitary, we instead find the closest unitary to $\hat{\mathcal{F}}_{m}$ by applying the SVD and discarding the singular values. We call this unitary $U_{\operatorname{new},m}$. However, replacing $U_{m}$ with $U_{\operatorname{new},m}$ is a strong update. Instead, we use a learning rate to smoothly interpolate between the two extremes. In this work, we use a linear learning rate
$U^{\prime}_{m}=(1-\beta)\,U_{m}+\beta\,U_{\operatorname{new},m},$ (43)
where $\beta\in(0,1)$ is the learning rate appropriately chosen beforehand. We note here that, as defined above, $U^{\prime}_{m}$ is not unitary. To mitigate this, we perform an SVD and discard the singular values again.
Algorithm 2 Optimizing Decomposition (OD)
0: Input: Initial quantum circuit $\prod_{m=1}^{M}U_{m,0}$ Number of sweeps $T$ Target fidelity $\hat{f}$ Learning rate $r\in[0,1]$
0:Output: Optimized quantum circuit $\prod_{m=1}^{M}U_{m}$
1: Set $t=1$.
2:while $t\leq T$ and $|\langle 0^{\otimes N}|\prod_{m=M}^{1}U_{m}^{\dagger}|\psi_{\chi_{\max}}\rangle|^{2}<\hat{f}$ do
3:  for $m$ in $1\ldots M$ do
4:   Calculate $\hat{\mathcal{F}}_{m}$.
5:   Split using SVD $\hat{\mathcal{F}}=\mathcal{U}\mathcal{S}\mathcal{V}^{\dagger}$.
6:   Define new unitary $U_{\text{new,}m}=\mathcal{U}\mathcal{V}^{\dagger}$
7:   Apply learning rate to get new $U_{m}$.
8:  end for
9:  Increment $t=t+1$.
10:end while